OpenStudy (anonymous):

prove that 2^n>n for every n with mathematical induction

OpenStudy (zehanz):

First, you must have a value of n for which the inequality is obviously true. Can you name that n?

OpenStudy (anonymous):

the initial value of n n=1 => 2^1>1=>2>1 so its true for the initial value of n

OpenStudy (anonymous):

then we suppose that the property is true for n=k so 2^k>k

OpenStudy (zehanz):

This is going really well!

OpenStudy (anonymous):

now i am stuck in the first step for n=k+1 2^(k+1)>(k+1)

OpenStudy (zehanz):

OK, can you write 2^(k+1) in another way?

OpenStudy (anonymous):

Consider: \[ 2^k > k \implies 2^k+2^k > k + k \]

OpenStudy (anonymous):

yes 2^(k+1)=2x2^k

OpenStudy (zehanz):

Hey @wio, why are you giving it away?

OpenStudy (anonymous):

I haven't given it away.

OpenStudy (zehanz):

If you say so... Ok, so now you have 2^(k+1) = 2 x 2^k. You know something about 2^k...

OpenStudy (anonymous):

2^k>k

OpenStudy (zehanz):

Yes, so 2^(k+1)=2x2^k > 2k

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now i have to prove that 2k>k+1?

OpenStudy (zehanz):

2k=k+k, so 2^(k+1)=2x2^k > 2k = k + k. What do you have to prove in this step?

OpenStudy (anonymous):

that k+k>k+1?

OpenStudy (zehanz):

You have to prove that 2^(k+1)> k+1. You have proven: 2^(k+1)> k+k. I think we're done, because k+k > k+1, if k > 1. This is OK, because the case k=1 is not interesting (it was the initial step) So you have proven the inequality for n=k+1. Because the inequality is true for n=1, it is now true for n=1+1=2, for n=2+1=3 and so on.

OpenStudy (anonymous):

No, \(k \not > 1\). \(k\geq 1\). If you ignore \(k=1\) then the \(k=2\) case needs to be proven independently, because you fail to prove \(p(2)\) and thus \(p(2)\implies p(3)\) is not sufficient to prove \(p(3)\) and so on.

OpenStudy (zehanz):

@wio is right. Taking k=1 into account, we have proven: 2^(k+1)> k+k ≥ k+1. So if you compare the leftmost term with the rightmost you have 2^(k+1) > k+1. Now we're done.